A bicycle rider travels 250m north and 160 m northwest. What is the riders displacement? What distance does the rider cover?
1
Expert's answer
2019-04-16T09:42:25-0400

The rider's displacement:


"\vec{s}=\vec{s_1}+\vec{s_2}"

"\vec{ s_1}=(0, 250)m"

"\vec{ s_2}=(160 \cos{45}, 160 \sin{45})m"

"\vec{s}=(0, 250)+(160 \cos{45}, 160 \sin{45})=(113.1, 363.1)m"

The magnitude is


"s=\sqrt{113.1^2+363.1^2}=380 m"

The direction is


"a=\arctan{\frac{113.1}{363.1}}=17.3^\circ"

west of north.

The distance is


"d=s_1+s_2=250+160=410 m."